OpenStudy (candy13106):

solve the linear system by using substitution. 3x+3y=-18 4x-y=-14

OpenStudy (campbell_st):

take the 2nd equation and make y the subject. so 4x + 14 = y and substitute it into the 1st equation 3x + 3(4x + 14) = -18 or 3x + 12x + 42 = -18 now solve for x

OpenStudy (candy13106):

idk how

OpenStudy (campbell_st):

oops ... collect like terms 15x + 42 = -18 subtract 42 from both sides of the equation then divide by 15

OpenStudy (candy13106):

thanks

OpenStudy (campbell_st):

when you get a value for x, substitute it into either equation to find y

OpenStudy (candy13106):

oh ok